OpenStudy (anonymous):

What is the most probable next item in the pattern 100, 81, 64, 49, ... ?

jimthompson5910 (jim_thompson5910):

Notice how 49 = 7^2 = 7*7, 64 = 8^2 = 8*8, etc etc

OpenStudy (anonymous):

Aren't you seeing a square pattern there ??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

can you say what will be the next term ??

OpenStudy (anonymous):

here you have a series that follows a pattern such as \[[x(n)] ^{2}\] Now as we proceed from 100 to 49in the series, these are the squares of 10,9,8,7 and so the next number will be the square of 6 i.e 36.

OpenStudy (anonymous):

100 is square of 10 81 is square of 9 64 is square of 8 49 is square of 7. .... is square of ... ??